# Bialgebras, and Lie monoid actions in Morse and Floer theory, I

### v1

@@ 2.4. Partial compactification
We have $\left|\underline{k}^{1}\sharp\underline{k}^{0}\right|=\left|\underline{k}^{1}\right|$, $n(\underline{k}^{1}\sharp\underline{k}^{0})=n(\underline{k}^{0})$, and this operation is associative. Given $\underline{k}$, there are $2^{v(\underline{k})}$ ways to split $\underline{k}$ as $\underline{k}^{1}\sharp\underline{k}^{0}$: intuitively, each vertex of $\underline{k}$ can go either in $\underline{k}^{0}$ or $\underline{k}^{1}$.
Definition 2.9.
Given a pair of multi-indices $(\underline{k},\underline{l})$, we call $n$-broken biforest multi-index for $(\underline{k},\underline{l})$ sequences $(\underline{\underline{k}},\underline{\underline{l}})$ of multi-indices
(2.34) $\underline{\underline{k}}$ $=(\underline{k}^{0},\ldots,\underline{k}^{n}),$
(2.35) $\underline{\underline{l}}$ $=(\underline{l}^{0},\ldots,\underline{l}^{n}),$
such that, for $0\leq f\leq n$, $\underline{k}^{f}$ and $\underline{l}^{f}$ are not simultaneously vertical, and such that
(2.36) $\underline{k}$ $=\underline{k}^{n}\sharp\cdots\sharp\underline{k}^{0},$
(2.37) $\underline{l}$ $=\underline{l}^{0}\sharp\cdots\sharp\underline{l}^{n}.$
As in Symplectic Field Theory [EGH00], we think of an $n$-broken biforest as forming a building, and $(\underline{k}^{f},\underline{l}^{f})$ is referred to as the biforest at floor $f$ (recall that ground floor is floor zero).
Some sequences of multi-indices are redundant: when facing a vertical forest, a tree breaking in subtrees can be represented “in levels” in many ways, see figure 9. Therefore we define an equivalence relation on the set of broken biforests multi-indices. In codimension one, these subtleties can essentially be ignored, as we will see in Section 2.5.
$\underline{\underline{k}}_{3}$
$\underline{\underline{k}}_{2}$
$\underline{\underline{k}}_{1}$
$\underline{\underline{k}}_{3}$
$\underline{\underline{k}}_{2}^{\prime}$
$\underline{\underline{k}}_{1}$
Figure 9. Two equivalent broken biforests.
Recall that ${K}_{k}$ compactifies to $\overline{K}_{k}$ by adding broken trees. A broken forest $\underline{\underline{k}}$ such that $k=\underline{k}^{n}\sharp\cdots\sharp\underline{k}^{0}$ determines a broken tree by associating a vertex to each entry $\geq 2$, and connect them by edges in the obvious way. Though, several broken forest multi-indices can represent the same broken tree.
Definition 2.10.
Assume as in Figure 9 that $(\underline{\underline{k}},\underline{\underline{l}})$ and $(\underline{\underline{k}}^{\prime},\underline{\underline{l}}^{\prime})$ are broken biforest multi-indices of length $n=n_{1}+n_{2}+n_{3}$ and $n^{\prime}=n_{1}+n_{2}^{\prime}+n_{3}$ respectively, and can be written as:
(2.38) $\underline{\underline{k}}$ $=(\underline{\underline{k}}_{1},\underline{\underline{k}}_{2},\underline{\underline{k}}_{3}),$
(2.39) $\underline{\underline{l}}$ $=(\underline{\underline{l}}_{1},\underline{\underline{l}}_{2},\underline{\underline{l}}_{3}),$
(2.40) $\underline{\underline{k}}^{\prime}$ $=(\underline{\underline{k}}_{1},\underline{\underline{k}}_{2}^{\prime},\underline{\underline{k}}_{3}),$
(2.41) $\underline{\underline{l}}^{\prime}$ $=(\underline{\underline{l}}_{1},\underline{\underline{l}}_{2}^{\prime},\underline{\underline{l}}_{3}),$
Assume that $\underline{\underline{l}}_{2}$ and $\underline{\underline{l}}_{2}^{\prime}$ are vertical broken forests, and that both $\underline{\underline{k}}_{2}$ and $\underline{\underline{k}}_{2}^{\prime}$ glue to the same multi-index $\tilde{\underline{k}}$. If $\underline{\underline{k}}_{2}$ and $\underline{\underline{k}}_{2}^{\prime}$ define the same face of $\overline{K}_{\tilde{\underline{k}}}$, then we say that $(\underline{\underline{k}},\underline{\underline{l}})$ and $(\underline{\underline{k}}^{\prime},\underline{\underline{l}}^{\prime})$ are equivalent. Likewise exchanging the roles of $k$ and $l$.
This generates an equivalence relation on the set of broken biforest multi-indices, and we denote $BF(\underline{k},\underline{l})$ its quotient. It will serve as indexing set for the boundary strata of $\overline{K}^{\underline{k}}_{\underline{l}}$.
For $[\underline{\underline{k}},\underline{\underline{l}}]\in BF(\underline{k},\underline{l})$, let the corresponding space of broken biforests
(2.42) $K^{\underline{\underline{k}}}_{\underline{\underline{l}}}=K^{\underline{k}^{0}}_{\underline{l}^{0}}\times\cdots\times K^{\underline{k}^{n}}_{\underline{l}^{n}}$
Notice that this space is independent555as an unoriented vector space. of the representative of $[\underline{\underline{k}},\underline{\underline{l}}]$. Its codimension (as a stratum of $\overline{K}^{\underline{k}}_{\underline{l}}$) is defined as
(2.43) $c(\underline{\underline{k}},\underline{\underline{l}})$ $=\dim{K}^{\underline{k}}_{\underline{l}}-\dim K^{\underline{\underline{k}}}_{\underline{\underline{l}}}$
(2.44) $=\sum_{i=0}^{n}c(\underline{k}^{i},\underline{l}^{i})-c(\underline{k},\underline{l}).$
This quantity is bounded above and below, with $n$ the length of $(\underline{\underline{k}},\underline{\underline{l}})$:
(2.45) $n\leq c(\underline{\underline{k}},\underline{\underline{l}})\leq\dim{K}^{\underline{k}}_{\underline{l}}.$
By denoting $BF_{c}(\underline{k},\underline{l})\subset BF(\underline{k},\underline{l})$ the subset of those $[\underline{\underline{k}},\underline{\underline{l}}]$ for which $c(\underline{\underline{k}},\underline{\underline{l}})=c$, we have a partition
(2.46) $BF(\underline{k},\underline{l})=BF_{0}(\underline{k},\underline{l})\cup BF_{1}(\underline{k},\underline{l})\cup\cdots\cup BF_{\dim{K}^{\underline{k}}_{\underline{l}}}(\underline{k},\underline{l}).$
Likewise, for the boundary strata of $\overline{J}^{\underline{k}}_{\underline{l}}$:
Definition 2.11.
Say that $(\underline{\underline{k}},\underline{\underline{l}};i)$ is a 1-grafted broken biforest multi-index for $(\underline{k},\underline{l})$ if $i$ is an integer $0\leq i\leq n$, corresponding to the grafting position, and $(\underline{\underline{k}},\underline{\underline{l}})$ is a broken biforest multi-index as above, except that $\underline{k}^{i}$ and $\underline{l}^{i}$ are allowed to be simultaneously vertical. To such a triple we will associate the corresponding space of grafted broken biforests:
(2.47) $J^{\underline{\underline{k}}}_{\underline{\underline{l}};i}=K^{\underline{k}^{0}}_{\underline{l}^{0}}\times\cdots\times K^{\underline{k}^{i-1}}_{\underline{l}^{i-1}}\times J^{\underline{k}^{i}}_{\underline{l}^{i}}\times K^{\underline{k}^{i+1}}_{\underline{l}^{i+1}}\times\cdots\times K^{\underline{k}^{n}}_{\underline{l}^{n}}.$
Just as for $K^{\underline{\underline{k}}}_{\underline{\underline{l}}}$, some triples are equivalent. The equivalence relation is defined analogously, except that we require furthermore that $i$ (resp. $i^{\prime}$) does not meet $(\underline{\underline{k}}_{2},\underline{\underline{l}}_{2})$ (resp. $(\underline{\underline{k}}_{2}^{\prime},\underline{\underline{l}}_{2}^{\prime})$). If $i$ meets $(\underline{\underline{k}}_{1},\underline{\underline{l}}_{1})$, then we want $i^{\prime}=i$, and if $i$ meets $(\underline{\underline{k}}_{3},\underline{\underline{l}}_{3})$, then we want $i^{\prime}$ to meet $(\underline{\underline{k}}_{3},\underline{\underline{l}}_{3})$ at the same position, i.e. $i^{\prime}-n_{2}^{\prime}=i-n_{2}$.
We denote $GBF(\underline{k},\underline{l})$ the quotient. It will serve as indexing set for the boundary strata of $\overline{J}^{\underline{k}}_{\underline{l}}$.
The codimension of $J^{\underline{\underline{k}}}_{\underline{\underline{l}};i}$ (as a stratum of $\overline{J}^{\underline{k}}_{\underline{l}}$) is defined as
(2.48) $c(\underline{\underline{k}},\underline{\underline{l}};i)$ $=\dim{J}^{\underline{k}}_{\underline{l}}-\dim J^{\underline{\underline{k}}}_{\underline{\underline{l}};i}$
(2.49) $=\sum_{j\neq i}c(\underline{k}^{j},\underline{l}^{j}).$
This quantity is bounded above and below:
(2.50) $n\leq c(\underline{\underline{k}},\underline{\underline{l}};j)\leq\dim{J}^{\underline{k}}_{\underline{l}}.$
By denoting $GBF_{c}(\underline{k},\underline{l})\subset GBF(\underline{k},\underline{l})$ the subset of those $(\underline{\underline{k}},\underline{\underline{l}};i)$ for which $c(\underline{\underline{k}},\underline{\underline{l}};i)=c$, we have a partition
(2.51) $GBF(\underline{k},\underline{l})=GBF_{0}(\underline{k},\underline{l})\cup GBF_{1}(\underline{k},\underline{l})\cup\cdots\cup GBF_{\dim{J}^{\underline{k}}_{\underline{l}}}(\underline{k},\underline{l}).$
As sets, $\overline{K}^{\underline{k}}_{\underline{l}}$ and $\overline{J}^{\underline{k}}_{\underline{l}}$ are defined as the unions:
(2.52) $\overline{K}^{\underline{k}}_{\underline{l}}$ $=\bigcup_{(\underline{\underline{k}},\underline{\underline{l}})\in BF(\underline{k},\underline{l})}{K}^{\underline{\underline{k}}}_{\underline{\underline{l}}},$
(2.53) $\overline{J}^{\underline{k}}_{\underline{l}}$ $=\bigcup_{(\underline{\underline{k}},\underline{\underline{l}};i)\in GBF(\underline{k},\underline{l})}{J}^{\underline{\underline{k}}}_{\underline{\underline{l}};i}.$
As topological spaces (manifolds with generalized corners) we glue all the stratum inductively via glueing maps.
If $(\underline{\underline{k}},\underline{\underline{l}})\in BF_{1}(\underline{k},\underline{l})$, then one can assume that $n=1$, and one can define a glueing map
(2.54) $g^{\underline{\underline{k}}}_{\underline{\underline{l}}}\colon[0,+\infty)\times{K}^{\underline{\underline{k}}}_{\underline{\underline{l}}}\to\overline{K}^{\underline{k}}_{\underline{l}},$
with $h\in[0,+\infty)$, by picking the biforest representatives $(\underline{U}^{0},\underline{D}^{0})$ and $(\underline{U}^{1},\underline{D}^{1})$ so that the relative height between the two vertices to be glued equals $h$, and taking the glued forest as in Figure 10. Notice that {+}×$\left\{+\infty\right\}\times{K}^{\underline{\underline{k}}}_{\underline{\underline{l}}}$ matches the orientation of ${K}^{\underline{\underline{k}}}_{\underline{\underline{l}}}$, as a boundary of $[0,+\infty]\times{K}^{\underline{\underline{k}}}_{\underline{\underline{l}}}$.
@@
Figure 10. Glueing two biforests.
One can then glue along the $g^{\underline{\underline{k}}}_{\underline{\underline{l}}}$:
@@
(2.55) $\left(\overline{K}^{\underline{k}}_{\underline{l}}\right)_{\leq 1}=\left({K}^{\underline{k}}_{\underline{l}}\cup\bigcup_{(\underline{\underline{k}},\underline{\underline{l}})\in BF_{1}(\underline{k},\underline{l})}[0,+\infty]\times{K}^{\underline{\underline{k}}}_{\underline{\underline{l}}}\right){/_{\sim}}.$
Likewise, if $(\underline{\underline{k}},\underline{\underline{l}};i)\in GBF_{1}(\underline{k},\underline{l})$, then one can assume that $n=1$, and one can define a glueing map
(2.56) $g^{\underline{\underline{k}}}_{\underline{\underline{l}};i}\colon[0,+\infty)\times{J}^{\underline{\underline{k}}}_{\underline{\underline{l}};i}\to\overline{J}^{\underline{k}}_{\underline{l}},$
and glue the boundary along it:
@@
(2.57) $\left(\overline{J}^{\underline{k}}_{\underline{l}}\right)_{\leq 1}=\left({J}^{\underline{k}}_{\underline{l}}\cup\bigcup_{(\underline{\underline{k}},\underline{\underline{l}};i)\in GBF_{1}(\underline{k},\underline{l})}[0,+\infty]\times{J}^{\underline{\underline{k}}}_{\underline{\underline{l}};i}\right){/_{\sim}}.$
2.5. Explicit description of codimension zero and one strata
We now describe $BF_{c}(\underline{k},\underline{l})$ and $GBF_{c}(\underline{k},\underline{l})$ more explicitly, when $c=0,1$.
$BF_{0}(\underline{k},\underline{l})$ : A priori, $n$ can take any values between 00 and $c(\underline{k},\underline{l})-1$, since if $c(\underline{k},\underline{l})>1$ nonvertical trees can be placed in different levels. However, these are all equivalent to $n=0$, $(\underline{\underline{k}},\underline{\underline{l}})=(\underline{k},\underline{l})$. Therefore one can write $BF_{0}(\underline{k},\underline{l})=\left\{(\underline{k},\underline{l})\right\}$.
$BF_{1}(\underline{k},\underline{l})$ : In the generic situation when both $(\underline{k},\underline{l})$ are nonvertical, then one must have $n=1$ and the equivalence relation is vacuous. In this case $BF_{1}(\underline{k},\underline{l})$ consists in broken biforests $(\underline{\underline{k}},\underline{\underline{l}})=(\underline{k}^{0},\underline{k}^{1};\underline{l}^{0},\underline{l}^{1})$ such that $\underline{k}=\underline{k}^{1}\sharp\underline{k}^{0}$, $\underline{l}=\underline{l}^{0}\sharp\underline{l}^{1}$, and for $f=0,1$, $c(\underline{k}^{f},\underline{l}^{f})=1$, i.e. if either $\underline{k}^{f}$ or $\underline{l}^{f}$ is vertical, then the other contains exactly one nonvertical tree.
In the case when one forest is vertical, then just as for $BF_{0}(\underline{k},\underline{l})$, $n$ might be as large as $c(\underline{k},\underline{l})$, though one can always merge floors and assume $n=1$. The equivalence relation is nontrivial here: the two trees facing each other are fixed, but the other “free” trees can move from one floor to the other, as in Figure 11. We will kill this ambiguity by assuming, by convention, that all these free trees must be in the lowest floor, i.e. floor 0. That means that floor $1$ only has one nonvertical tree, and it faces one nonvertical tree of floor 0. Equivalently, $c(\underline{k}^{1},\underline{l}^{1})=1$.
Figure 11. Two broken forests leading to equivalent broken biforests when facing a vertical descending forest.
$GBF_{0}(\underline{k},\underline{l})$ : One must have $n=0$, and the equivalence relation is vacuous.
$GBF_{1}(\underline{k},\underline{l})$ : Again, $n$ might be as large as $c(\underline{k},\underline{l})$, but can be assumed to equal 1. Then the equivalence relation becomes vacuous, and
(2.58) $GBF_{1}(\underline{k},\underline{l})=GBF_{1}^{f=0}(\underline{k},\underline{l})\cup GBF_{1}^{f=0}(\underline{k},\underline{l}),$
where $GBF_{1}^{f=0}(\underline{k},\underline{l})$ corresponds to $(\underline{\underline{k}},\underline{\underline{l}})$ such that $c(\underline{k}^{1},\underline{l}^{1})=1$ with no constraints on $(\underline{k}^{0},\underline{l}^{0})$, and vice versa for $GBF_{1}^{f=1}(\underline{k},\underline{l})$.
To summarize, the indexing sets that will appear in the coherence relations are:
@@ 2.5. Explicit description of codimension zero and one strata
(2.59) $BF_{1}(\underline{k},\underline{l})$ =$=\left\{(\underline{k}^{0},\underline{l}^{0},\underline{k}^{1},\underline{l}^{1})\,\left|\,\begin{aligned} &\underline{k}=\underline{k}^{1}\sharp\underline{k}^{0},\,\underline{l}=\underline{l}^{0}\sharp\underline{l}^{1},\\ &c(\underline{k}^{0},\underline{l}^{0})=c(\underline{k},\underline{l}),\\ &c(\underline{k}^{1},\underline{l}^{1})=1\end{aligned}\right.\right\}.$
@@
(2.60) $GBF_{1}^{f=0}(\underline{k},\underline{l})$ =$=\left\{(\underline{k}^{0},\underline{l}^{0},\underline{k}^{1},\underline{l}^{1})\,\left|\,\begin{aligned} &\underline{k}=\underline{k}^{1}\sharp\underline{k}^{0},\,\underline{l}=\underline{l}^{0}\sharp\underline{l}^{1},\\ &c(\underline{k}^{1},\underline{l}^{1})=1\end{aligned}\right.\right\}.$
(2.61) $GBF_{1}^{f=1}(\underline{k},\underline{l})$ =$=\left\{(\underline{k}^{0},\underline{l}^{0},\underline{k}^{1},\underline{l}^{1})\,\left|\,\begin{aligned} &\underline{k}=\underline{k}^{1}\sharp\underline{k}^{0},\,\underline{l}=\underline{l}^{0}\sharp\underline{l}^{1},\\ &c(\underline{k}^{0},\underline{l}^{0})=1\end{aligned}\right.\right\}.$

### v2

@@ 2.4. Partial compactification
We have $\left|\underline{k}^{1}\sharp\underline{k}^{0}\right|=\left|\underline{k}^{1}\right|$, $n(\underline{k}^{1}\sharp\underline{k}^{0})=n(\underline{k}^{0})$, and this operation is associative. Moreover, one has a natural identification
(2.36) $\mathrm{Vert}(\underline{k}^{1}\sharp\underline{k}^{0})\simeq\mathrm{Vert}(\underline{k}^{1})\cup\mathrm{Vert}(\underline{k}^{0}).$
Indeed, one has $\mathrm{Vert}(\underline{k}^{1})\subset\mathrm{Vert}(\underline{k}^{1}\sharp\underline{k}^{0})$, and $\mathrm{Vert}(\underline{k}^{1}\sharp\underline{k}^{0})\setminus\mathrm{Vert}(\underline{k}^{1})$ is in an order-preserving one-to-one correspondence with $\mathrm{Vert}(\underline{k}^{0})$.
In particular, given $\underline{k}$, there are $2^{v(\underline{k})}$ ways to split $\underline{k}$ as $\underline{k}^{1}\sharp\underline{k}^{0}$, which corresponds to the number of ways of splitting $\mathrm{Vert}(\underline{k})$ into two subsets.
It follows from (2.36) and the identification (2.9) that one has a natural (unoriented, the sign change will be computed in Section 2.5) identification
(2.37) $J^{\underline{k}}_{\underline{l}}\simeq J^{\underline{k}^{0}}_{\underline{l}^{0}}\oplus J^{\underline{k}^{1}}_{\underline{l}^{1}}.$
Denote
(2.38) $v^{i}$ $=(1,\ldots,1)\in G^{\underline{k}^{i}}_{\underline{l}^{i}}\subset J^{\underline{k}^{i}}_{\underline{l}^{i}}\simeq\mathbb{R}^{v(\underline{k}^{i})+v(\underline{l}^{i})},\text{ and}$
(2.39) $v$ $=v^{0}+v^{1}=(1,\ldots,1)\in G^{\underline{k}}_{\underline{l}}\subset J^{\underline{k}}_{\underline{l}}.$
If $c(\underline{k}^{1},\underline{l}^{1})=1$, then one has $J^{\underline{k}^{1}}_{\underline{l}^{1}}\simeq\mathbb{R}v^{1}\oplus K^{\underline{k}^{1}}_{\underline{l}^{1}}$. Call the resulting identification the gluing map:
(2.40) $g^{\underline{k}^{0},\underline{k}^{1}}_{\underline{l}^{0},\underline{l}^{1};0}\colon\mathbb{R}\times J^{\underline{k}^{0}}_{\underline{l}^{0}}\times K^{\underline{k}^{1}}_{\underline{l}^{1}}$ $\to J^{\underline{k}}_{\underline{l}}$
(2.41) $(h,B_{0},B_{1})$ $\mapsto hv_{1}+B_{0}+B_{1}.$
Observe that when the parameter $h$ is sufficiently large, this corresponds to gluing the biforest of $K^{\underline{k}^{1}}_{\underline{l}^{1}}$ (shifted up by $h$) on top of the one of $J^{\underline{k}^{0}}_{\underline{l}^{0}}$, as in Figure 9.
@@
Figure 9. Gluing two biforests.
Likewise, if $c(\underline{k}^{0},\underline{l}^{0})=1$, then one has $J^{\underline{k}^{0}}_{\underline{l}^{0}}\simeq\mathbb{R}v^{0}\oplus K^{\underline{k}^{0}}_{\underline{l}^{0}}$. Identify $\mathbb{R}\simeq\mathbb{R}(-v^{0})$, and call the resulting identification the gluing map:
(2.42) $g^{\underline{k}^{0},\underline{k}^{1}}_{\underline{l}^{0},\underline{l}^{1};1}\colon\mathbb{R}\times K^{\underline{k}^{0}}_{\underline{l}^{0}}\times J^{\underline{k}^{1}}_{\underline{l}^{1}}$ $\to J^{\underline{k}}_{\underline{l}}$
(2.43) $(h,B_{0},B_{1})$ $\mapsto-hv_{0}+B_{0}+B_{1}.$
This corresponds to gluing the biforest of $K^{\underline{k}^{0}}_{\underline{l}^{0}}$, shifted down by $h\in\mathbb{R}$, below the one of $J^{\underline{k}^{1}}_{\underline{l}^{1}}$.
Assume now that $c(\underline{k},\underline{l})=c(\underline{k}^{0},\underline{l}^{0})=c(\underline{k}^{1},\underline{l}^{1})=1$, then let the gluing map
(2.44) $g^{\underline{k}^{0},\underline{k}^{1}}_{\underline{l}^{0},\underline{l}^{1}}\colon\mathbb{R}\times K^{\underline{k}^{0}}_{\underline{l}^{0}}\times K^{\underline{k}^{1}}_{\underline{l}^{1}}\to K^{\underline{k}}_{\underline{l}}.$
be obtained by composing the map
(2.45) ×$\mathbb{R}\times K^{\underline{k}^{0}}_{\underline{l}^{0}}\times K^{\underline{k}^{1}}_{\underline{l}^{1}}$ $\to J^{\underline{k}}_{\underline{l}},$
(2.46) $(h,B_{0},B_{1})$ $\mapsto hv_{1}+B_{0}+B_{1},$
with the projection $J^{\underline{k}}_{\underline{l}}\to K^{\underline{k}}_{\underline{l}}$. Then define the partial compactifications:
@@
(2.47) $\left(\overline{K}^{\underline{k}}_{\underline{l}}\right)_{\leq 1}$ $=\left({K}^{\underline{k}}_{\underline{l}}\cup\bigcup_{\begin{subarray}{c}\underline{k}=\underline{k}^{1}\sharp\underline{k}^{0}\\ \underline{l}=\underline{l}^{0}\sharp\underline{l}^{1}\end{subarray}}(-\infty,+\infty]\times{K}^{\underline{k}^{0}}_{\underline{l}^{0}}\times{K}^{\underline{k}^{1}}_{\underline{l}^{1}}\right)/\sim$
@@
(2.48) $\left(\overline{J}^{\underline{k}}_{\underline{l}}\right)_{\leq 1}$ $=\left({J}^{\underline{k}}_{\underline{l}}\cup\bigcup_{\begin{subarray}{c}\underline{k}=\underline{k}^{1}\sharp\underline{k}^{0}\\ \underline{l}=\underline{l}^{0}\sharp\underline{l}^{1}\end{subarray}}(-\infty,+\infty]\times\left({J}^{\underline{k}^{0}}_{\underline{l}^{0}}\times{K}^{\underline{k}^{1}}_{\underline{l}^{1}}\cup{K}^{\underline{k}^{0}}_{\underline{l}^{0}}\times{J}^{\underline{k}^{1}}_{\underline{l}^{1}}\right)\right)/\sim,$
Where $\sim$ identifies points with their image under the gluing maps. These are manifolds with boundary:
@@
(2.49) $\partial\left(\overline{K}^{\underline{k}}_{\underline{l}}\right)_{\leq 1}$ $=\bigcup_{\begin{subarray}{c}\underline{k}=\underline{k}^{1}\sharp\underline{k}^{0}\\ \underline{l}=\underline{l}^{0}\sharp\underline{l}^{1}\end{subarray}}(-1)^{\rho}\cdot{K}^{\underline{k}^{0}}_{\underline{l}^{0}}\times{K}^{\underline{k}^{1}}_{\underline{l}^{1}},$
@@
(2.50) $\partial\left(\overline{J}^{\underline{k}}_{\underline{l}}\right)_{\leq 1}$ $=\bigcup_{\begin{subarray}{c}\underline{k}=\underline{k}^{1}\sharp\underline{k}^{0}\\ \underline{l}=\underline{l}^{0}\sharp\underline{l}^{1}\end{subarray}}\left((-1)^{\rho_{0}}\cdot{J}^{\underline{k}^{0}}_{\underline{l}^{0}}\times{K}^{\underline{k}^{1}}_{\underline{l}^{1}}\cup(-1)^{\rho_{1}}\cdot{K}^{\underline{k}^{0}}_{\underline{l}^{0}}\times{J}^{\underline{k}^{1}}_{\underline{l}^{1}}\right),$
where ${\rho=\rho^{\underline{k}^{0},\underline{k}^{1}}_{\underline{l}^{0},\underline{l}^{1}}},{\rho_{0}=\rho^{\underline{k}^{0},\underline{k}^{1}}_{\underline{l}^{0},\underline{l}^{1};0}},{\rho_{1}=\rho^{\underline{k}^{0},\underline{k}^{1}}_{\underline{l}^{0},\underline{l}^{1};1}}\in\mathbb{Z}_{2}$ stand for the orientations of the gluing maps ${g^{\underline{k}^{0},\underline{k}^{1}}_{\underline{l}^{0},\underline{l}^{1}}},{g^{\underline{k}^{0},\underline{k}^{1}}_{\underline{l}^{0},\underline{l}^{1};0}},{g^{\underline{k}^{0},\underline{k}^{1}}_{\underline{l}^{0},\underline{l}^{1};1}}$, which will appear in the coherence relations of $f$-bialgebras ($R^{\underline{k}}_{\underline{l}}$) (resp. of morphisms of $f$-algebras ($M^{\underline{k}}_{\underline{l}}$)); and which we are going to determine now.